The line that has a slope 1/2 and contains (5,6)
Answer by venugopalramana(3286)

The line that has a slope M=1/2 and contains P=(X1,Y1)=(5,6)
FORMULA IS
Y-Y1=M(X-X1)
Y-6=0.5(X-5)
2Y-12=X-5
2Y-X-7=0